Let V be the set of all positive real numbers; define + by u+v = uv (+ is ordinary multiplication) and define • by c•v = . Prove that V is a vector space.



1
Expert's answer
2022-03-15T14:17:18-0400

"Solution: If~u,v \in V~and ~c \in R,we~have~u+v=uv \in V,c.v=v^c \in V, So~ V~is~
\\closed~under~addition~and~scalar~multiplication.
\\Now ~we ~prove ~ remaining ~axioms:
\\1) We~have
\\~~~~~~~~u+v=uv(\because addition)
\\~~~~~~~~~~~~~~~~~~~=vu(\because multiplication~cummutative~in~R)
\\~~~~~~~~~~~~~~~~~~~=v+u (\because addition)
\\2) Note ~that
\\~~~~~~~~~~(u+v)+w=(uv)+w(\because addition)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=(uv)w(\because multiplication~cummutative~in~R)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=u(vw)(\because multiplication~ associative ~in~R)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=u+(vw) (\because addition)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=u+(v+w) (\because addition)
\\3)Observe~that~1\in V~and~1+u=1u=u
\\4) For~any~u \in V, note~that~u^{-1} \in V~so~that~u+u^{-1}=uu^{-1}=1
\\5) Note~that
\\~~~~~~~~~~~~~~~~~~~~~1.u=u^1=u 
\\6)Let~c,d\in R.We~have~
\\~~~~~~~~~~~~~~~~~~~~~(cd).u=u^{cd}~~~~~(\because scalar ~multiplication) 
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=(u^d)^c ~~~(\because~exponents ~in~R)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=c.(u^d)~~~~~(\because scalar ~multiplication)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=c.(d.u)~~~~~(\because scalar ~multiplication)
\\7)Note~that
\\~~~~~~~~~~~~~~~~~~~~~~c.(u+v)=c.(uv)~~~(\because addition)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=(uv)^c~~(\because scalar ~multiplication)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=u^cv^c ~~~(\because~exponents ~in~R)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=(u^c)+(v^c)~~(\because scalar ~multiplication)
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=(c.u)+(c.v)~~(\because addition)
\\8)We~have
\\~~~~~~~~(c+d).u=u^{c+d} (\because scalar~multiplication)
\\~~~~~~~~~~~~~~~~~~~~~~~~=u^c.u^d~~~~(\because exponents ~in~R)
\\~~~~~~~~~~~~~~~~~~~~~~~~=(u^c)+(u^d)~~~~(\because~addition)
\\~~~~~~~~~~~~~~~~~~~~~~~~=(c.u)+(d.u)~~(\because~Scalar~multiplication)
\\All~axioms ~are~satisfied.Hence~V~is ~ a~vector~space."